OpenStudy (anonymous):

If each H2molecule produces2e -how many kilograms of hydrogen would be required to generate the electricity needed for a typical house?

OpenStudy (anonymous):

Do you have a stated amount for the average amount of electric a typical house uses? I doubt they'd have asked this without more information.

OpenStudy (anonymous):

Also, what is the 'weight' of an electron according to your book/homework?

OpenStudy (anonymous):

Ok now the thing is we have got 2e from each hydrogen

OpenStudy (anonymous):

@Ishaan: two from each diatomic molecule! Not just atom. Wanted to point that out.

OpenStudy (anonymous):

e charge is something about 1.6 x 10^-19 coulombs

OpenStudy (anonymous):

Find out how much it takes first. Then use a constant to know how many electrons you need (there is a coulomb amount/mol), then use the stoichometric coefficients (i'm assuming a 2) since: \[H_2 \rightarrow 2H^+2e^-\] Then use molar mass of H2 (2.002) to convert to grams from moles

OpenStudy (anonymous):

Awesome. Now we need the book's stated mass and the average household usage.

OpenStudy (anonymous):

but the thing is it's not current it is the power or voltage needed for electricity

OpenStudy (anonymous):

96.485 kJ per volt gram equivalent Thats for electrons. Its Faraday's constant.

OpenStudy (anonymous):

Pumpkin: Could you tell us everything the book provided?

OpenStudy (anonymous):

Awesome. So I guess that just leaves us with needing the average electric usage.

OpenStudy (anonymous):

Oh! I think I just understood what Ishaan was saying

OpenStudy (anonymous):

We're looking for voltage, not kw/h?

OpenStudy (anonymous):

lets make it 240 w it in india and in us it is 110

OpenStudy (anonymous):

Wouldn't the book most likely be american standard? I'm assuming that it's for an american class, seeing as it's in English.

OpenStudy (anonymous):

Now P= i^2 r r resistance and i current v = ir v for voltage i = v/r

OpenStudy (anonymous):

and lets assume we have copper wire for house electric wiring

OpenStudy (anonymous):

thats makes p= v^2/r

OpenStudy (anonymous):

yeah so american standard p =110 right?

OpenStudy (anonymous):

ah we can't solve it this way we need diameter and length of wire too where did she get this?

OpenStudy (anonymous):

220 vs 110. Sounds right to me.

OpenStudy (anonymous):

I am not very good with physics, but I'm getting the impression that pumpkin hasn't provided ALL the information that was given. :(

OpenStudy (anonymous):

yeah your right